# CLE-SH: Comprehensive Literal Explanation package for SHapley values by statistical validity

### v1

@@ Abstract
Recently, SHapley Additive exPlanations (SHAP) has been widely utilized in various research domains. This is particularly evident in medical applications, where SHAP analysis serves as a crucial tool for identifying biomarkers and assisting in result validation. However, despite its frequent usage, SHAP is often not applied in a manner that maximizes its potential contributions. A review of recent papers employing SHAP reveals that many studies subjectively select a limited number of features as ’important’ and analyze SHAP values by approximately observing plots without assessing statistical significance. Such superficial application may hinder meaningful contributions to the applied fields. To address this, we propose a library package designed to simplify the interpretation of SHAP values. By simply inputting the original data and SHAP values, our library provides: 1) the number of important features to analyze, 2) the pattern of each feature via univariate analysis, and 3) the interaction between features. All information is extracted based on its statistical significance and presented in simple, comprehensible sentences, enabling users of all levels to understand the interpretations. We hope this library fosters a comprehensive understanding of statistically valid SHAP results.
Keywords SHapley Additive exPlanations  $\cdot$ Statistical Validity  $\cdot$ Visualization
1 Introduction
1.1 Background
@@
Between April 1st, 2024, and April 7th, 2024, by searching for (SHAP[Title/Abstract]) AND (Machine Learning[Title/Abstract]) on the PubMed site, 38 papers that use machine learning and SHAP together were published. We analyze how papers in the biomedical fields utilize SHAP in Figure1 and Appendix Table 1. Among 33 accessible manuscripts, 84.8% of the papers include SHAP summary plots, which summarize the SHAP value distribution for each feature by its rank [2, 15, 52, 14, 18, 28, 41, 25, 45, 53, 1, 46, 48, 40, 20, 31, 39, 19, 43, 17, 49, 33, 26, 34, 38, 8, 24, 10, 7, 51, 35, 16, 11, 32, 6, 27, 29, 37].
Figure 1: Investigation on the SHAP analysis usage in recent papers in medical fields
1.2 Problem Definition
@@
1.3 Proposed Approach
@@
We propose a Python library package that allows any user to easily access and apply proper statistical methods to their SHAP values, generating a comprehensive literal report PDF file that even non-experts in data science can easily understand. Meanwhile, we allow any user to utilize this package as they intend by setting hyperparameters.
1) Statistical tests to analyze SHAP values: In Sections 2 to 5, we introduce a procedure to analyze SHAP values with statistical significance for any type of tabular dataset. First, we define the number of important features to focus on, which involves feature selection by SHAP values (Section 2). Next, we determine the data type of each important feature, whether it is binary, discrete, or continuous (Section 3). Considering the type of each feature, we suggest univariate analysis to analyze each feature individually (Section 4). Finally, we design an interactive analysis to filter out statistically significant results from SHAP dependence plots (Section 5).
2) Summary report generation: In Section 6, we summarize the statistically significant results into a report intended to be comprehensive for non-experts in data analysis. To achieve this, instead of only displaying plots, we also generate literal sentences that explain the significant patterns within the plots.
1.4 How to use the library
The description and installation guide for the CLE-SH package is accessible at: xxx.xxx.xxx. After installation, importing the package with import clesh will enable the analysis by executing clesh(features, SHAP values, label name) as in Figure2. Here, features indicates the dataframe containing the feature dataset used to train/test the machine learning model and calculate SHAP values. SHAP values are SHAP values which match to features. label name is required to generate more comprehensive literal sentences.
1.5 Examples of the application
To validate the generalizability and reproducibility, we test our package using examples from five different datasets from various biomedical areas. These datasets vary in the number of samples, features, feature types, and obtainable performance (Table 1, Appendix 1) [4, 36, 13, 3, 23]. We collect SHAP values by test features in each 5-fold cross-validation, ensuring that every data sample was included in the results. The XGBoost algorithm, a widely used tree-based ensemble method popular in biomedical research, is used for the machine learning model (Appendix 2). We use the area under the ROC curve (AUC) to calculate the performance for each dataset. Details of our testing environment are provided in Appendix 3.
@@ 1.5 Examples of the application
Table 1: Data Characteristics. MS*=Metabolic Syndrome, BC*=Breast Cancer, HF*=Heart Failure, DR*=Diabetic Retinopathy, IBD*=Inflammatory Bowel Disease
@@
2 Number of important features
@@
As statistical significance is a key factor in this study, we compare the average absolute SHAP values between features. As shown in Figure3.a, we calculate the statistical significance of difference between adjacently ranked features. When both groups of SHAP values satisfy the normal distribution test by the Shapiro Wilk test, a paired t-test is used (parametric); otherwise, the Wilcoxon rank sum test is used (non-parametric). By cutting the group of features by where statistical significance occurs, we determine the number of important features according to the number of cuts, as shown in Figure3.b. The minimal number of important features (candidate_num_min, default=10) and the maximal number of important features (candidate_num_max, default=20) are set by hyperparameters, as having too many important features can be ineffective and too few can be less informative. Between candidate_num_min and candidate_num_max, we select the number that shows the highest difference from the next cut, as shown in Figure3.c. Statistical significance is determined p-value<0.05, which 0.05 can be changed by adjusting the hyperparameter p_feature_selection. In the clesh_result folder (see Figure2), the Figureshowing the number of important features according to the number of cuts, as in Figures 3.d and 3.e, will be saved.
3 Identification of feature data type
@@
4 Univariate analysis
@@
Table 2: Univariate analysis in each dataset. Linear, Quadratic, and Sigmoid in the first column imply the number of features that fit best to each function with statistical significance.
@@
4.1 Binary and Discrete Types
We provide box plots of SHAP values in each category, as shown in Figures 4.a and 4.b.
@@ 4.1 Binary and Discrete Types
4.2 Continuous Types
SHAP values of continuous features are difficult to interpret using conventional statistics, as the relationship can be very complex to interpret or feature values can be too sparse. To balance the amount of information and utility, we first provide scatter plots (SHAP dependence plots without interaction analysis). Then, we apply three different functions for interpretation: linear (Eq. 1), quadratic (Eq. 2), and sigmoid (Eq. 3). By setting feature values as the input variable and SHAP values as the output, each function is fitted, and its statistical significance is checked. As our goal is to determine which function best represents the impact of a feature, we define statistical significance by the p-value of the coefficient of the input (x), denoted as a. If more than one function type is statistically significant, the Root Mean Square Error from real data points is used to select the best function type. The selected function is displayed in the scatter plot. As shown in Table 2, the function type that fits the most is the sigmoid function, followed by quadratic and linear. This indicates that the relationship between feature values and SHAP values is usually non-linear, underscoring the need to analyze more than just the SHAP summary plot. An example of the linear function is illustrated in Figure4.d, the quadratic function in Figure4.e, and the sigmoid functions in Figures 4.f and 4.g.
[…]
@@ 4.2 Continuous Types
5 Interaction analysis
Interaction analysis is performed for important features in Section 2, namely, target features. To define the interaction feature of each feature, we use the built-in function of SHAP library, shap.utils.approximate_interactions().
@@ 5 Interaction analysis
1. Target: Binary/Discrete, Interaction: Binary/Discrete. In each category of the interaction feature, the SHAP values of the target feature between different categories of the target feature are compared. The statistical test follows the method described in Section 4.1. When a target feature is binary, a two-sample t-test (parametric) or Mann-Whitney U test (non-parametric) is applied. When a target feature is discrete with multiple categories, a one-way ANOVA (parametric) or Kruskal-Wallis test (non-parametric) is used, followed by Tukey’s HSD post-hoc analysis to identify statistically significant differences between categories.
2. Target: Binary/Discrete, Interaction: Continuous. When the interaction feature is continuous, analyzing by its range can be difficult to interpret. To simplify, we group interaction values into two categories: those smaller than the average and those larger than the average. This allows us to treat the interaction feature as a binary feature and apply the analysis described in Section 5.1.
In Figure5.a, an example of box plots in interaction analysis is illustrated (dataset: MS). The target feature ’Albuminuria’ shows a negative slope in its SHAP values, regardless of the interaction feature’s value (Triglycerides). However, in Figures 5.b and 5.c, we observe that the SHAP values, where ’Albuminuria’s feature value is 2, show a statistically significant difference from other SHAP values only when Triglycerides is above its average. This indicates that the impact of Albuminuria = 2 is more significant when triglycerides are above the average.
3. Target: Continuous, Interaction: Binary/Discrete. SHAP values of the target feature are grouped by the category of the interaction feature. Each group of SHAP values is then regressed by the target feature values using the function type selected in Section 4.2 (among Eq.1 ,2 or 3). Functions that are fitted with statistical significance will be displayed in the scatter plots as shown in Figure5.d. If only one group (one category of the interaction feature) shows statistical significance for the fitting, there will be no comparison between categories of the interaction feature. If more than one group shows statistical significance for the fitting, a comparison is made to determine whether the pattern depends on the category of the interaction feature.
@@
4. Target: Continuous, Interaction: Continuous. As in Section 5.3, samples are divided into two groups based on whether the interaction value is below or above its average. The analysis follows the same procedure as in Section 5.3, treating the interaction feature as binary.
6 Conclusion
To the best of the authors’ knowledge, this study is the first to develop a Python library that automatically performs statistical analysis on SHAP values. As illustrated in Figure1, the library automatically generates a report file summarizing all statistically significant patterns. Report PDF files for each dataset are included within the source code.
Proper recognition of the package’s limitations is essential for correct usage. Real-world features are highly interactive and complex, meaning that SHAP analysis applied to machine learning models cannot illustrate every inherent detail. The package presented in this paper summarizes statistically significant SHAP values, but the implications of these significant patterns should be interpreted by domain experts for each specific dataset. It is important to note that interactions between features are not represented by univariate analysis, and even interaction analysis may not fully elucidate the mechanisms underlying feature interactions. We hope that this paper contributes not by finalizing the biomarker discovery process, but by extracting meaningful insights from machine learning models that can guide more targeted research projects.
For future work, we suggest improvements in visualization and interface design. Additionally, we plan to expand the range of analyses in both univariate and interaction contexts. While the current version of the library does not pose a significant computational burden, as it only analyzes precomputed SHAP values, future enhancements might require consideration of strategies to reduce computational load as the library becomes more complex.

### v2

@@ Abstract
Recently, SHapley Additive exPlanations (SHAP) has been widely utilized in various research domains. This is particularly evident in application fields, where SHAP analysis serves as a crucial tool for identifying biomarkers and assisting in result validation. However, despite its frequent usage, SHAP is often not applied in a manner that maximizes its potential contributions. A review of recent papers employing SHAP reveals that many studies subjectively select a limited number of features as ’important’ and analyze SHAP values by approximately observing plots without assessing statistical significance. Such superficial application may hinder meaningful contributions to the applied fields. To address this, we propose a library package designed to simplify the interpretation of SHAP values. By simply inputting the original data and SHAP values, our library provides: 1) the number of important features to analyze, 2) the pattern of each feature via univariate analysis, and 3) the interaction between features. All information is extracted based on its statistical significance and presented in simple, comprehensible sentences, enabling users of all levels to understand the interpretations. We hope this library fosters a comprehensive understanding of statistically valid SHAP results.
I Introduction
I-A Background
@@
Between April 1st, 2024, and April 7th, 2024, by searching for (SHAP[Title/Abstract]) AND (Machine Learning[Title/Abstract]) on the PubMed site, 38 papers that use machine learning and SHAP together were published. We analyze how papers in the biomedical fields utilize SHAP in Figure1 and Appendix Table 1. Among 33 accessible manuscripts, 84.8% of the papers include SHAP summary plots, which summarize the SHAP value distribution for each feature by its rank [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
Figure 1: Investigation on the SHAP analysis usage in recent papers in medical fields. All figures were generated based on the search results available at the time of submission. One article (item 7 in Appendix Table 1) was retracted after submission. The figures and corresponding statistics were retained to preserve transparency of the initial dataset, but this retracted article had no influence on the study’s conclusions.
I-B Problem Definition
@@
I-C Related Works
With the increasing popularity of SHAP analysis, there has been a growing number of studies leveraging SHAP values for deeper and more meaningful analyses. In this section, we summarize recent studies that address at least one of the issues outlined in Subsection I-B.
I-C1 SHAP Values for Feature Selection
As SHAP values have gained attention, numerous attempts have been made to utilize them for feature selection. As shown in Figure 1, most approaches adapt traditional rank-based filtering methods to SHAP values. However, some studies propose entirely new methodologies for feature selection based on SHAP.
  • BorutaSHAP: Similar to the permutation importance approach, BorutaSHAP creates shadow features by randomly shuffling input features using TreeSHAP. It observes changes in performance across multiple iterations and calculates p-values to assess significance [49].
  • Shapicant: This method permutes the labels instead of features. It first calculates SHAP values for the actual labels and then computes null SHAP values for permuted labels. By comparing these, it calculates p-values to determine statistically significant differences [50].
  • PowerSHAP: PowerSHAP introduces noise into the dataset across multiple iterations and computes SHAP values. It evaluates the significance of SHAP values by comparing them to the added noise, identifying features with consistently high SHAP scores [51].
  • LLPowerSHAP: Building on PowerSHAP, this method assumes that informative features will have significantly higher SHAP values compared to noise. It incorporates Logistic Loss SHAP and utilizes unseen data to identify informative features with minimal noise among the selected feature set [52].
While these approaches suggest advanced feature selection methods with concrete algorithms, we propose a feature selection algorithm that also utilizes SHAP values. Our approach addresses the limitation of existing methods, which often require additional computational steps for feature selection. This requirement conflicts with our goal of generating a report file solely based on the input features and their SHAP values. In Section III, we introduce a much lighter version of SHAP-based feature selection that eliminates the need for any supplementary experiments beyond the provided SHAP values.
I-C2 Analysis on SHAP Values
There are numerous studies that perform additional analyses on SHAP values. Statistical results derived from linear and logistic regression was used to guide feature selection[53]. The SHAPley EXplanation Randomization Test (SHAP-XRT) utilizes the Conditional Randomization Test (CRT) to assess the statistical significance of SHAP values in binary classification tasks, demonstrating how SHAP values can be employed to evaluate feature importance[54].
In addition, several studies investigate the inherent instability of SHAP values. Similar to observations in other feature importance methods[55], research has analyzed how SHAP values fluctuate due to random sampling[56]. Furthermore, other studies evaluate the validity of SHAP scores by examining their behavior within uncertainty regions, identifying the actual tight range of SHAP scores[57].
In Section II-B, we provide a list of statistics incorporated into our Python library. While each method represents well-known basic statistical techniques, to the best of the author’s knowledge, no prior studies have applied such a wide range of statistical analyses to SHAP values.
I-C3 Simpler Explanation of SHAP Values
Despite the widespread use of SHAP values, their interpretation often remains challenging for non-experts and can be subject to individual bias. To address this, several studies attempt to simplify SHAP explanations using large language models (LLMs). For example, one study uses pre-trained LLMs to analyze sample-specific risk indicators, providing enhanced interpretability [58]. Another line of research employs LLMs to generate textual explanations for SHAP plots, making the visual outputs more accessible [59].
While there have been attempts to interpret ML models in a textual context, there is still no consensus on which information to prioritize among the entire set of SHAP values. We believe that a deterministic step to extract the ”important” information from SHAP values should precede the process of illustrating it in a textual manner.
I-D Proposed Approach
@@
The objective of this study is to develop a Python library, named CLE-SH, that automatically performs statistically validated SHAP analyses on biomedical tabular data and generates interpretable reports for both experts and non-experts. Rather than serving as a final step that replaces all analyses, CLE-SH is intended to be used in the early exploratory stage, helping researchers design a clear analytical direction from complex, information-rich datasets. To achieve this objective, we designed the library to integrate statistical validation, feature-type classification, univariate and interaction analyses, and report generation within a unified automated pipeline.
1) Statistical tests to analyze SHAP values: In Sections III,IV,V,VI, we introduce a procedure to analyze SHAP values with statistical significance for any type of tabular dataset. First, we define the number of important features to focus on, which involves feature selection by SHAP values (Section III). Next, we determine the data type of each important feature, whether it is binary, discrete, or continuous (Section IV). Considering the type of each feature, we suggest univariate analysis to analyze each feature individually (Section V). Finally, we design an interactive analysis to filter out statistically significant results from SHAP dependence plots (Section VI).
2) Summary report generation: We summarize the statistically significant results into a report intended to be comprehensive for non-experts in data analysis. To achieve this, instead of only displaying plots, we also generate literal sentences that explain the significant patterns within the plots (Appendix 6)
II Methods
II-A Datasets
To validate the generalizability and reproducibility, we test our package using examples from five different datasets from various biomedical areas. The data sources and label definitions for each dataset are described in Appendix 1, while detailed information on the number of samples, features, feature types, and obtainable performance is summarized in Table 1, demonstrating the diversity of the datasets [60, 61, 62, 63, 9]. We collect SHAP values by test features in each 5-fold cross-validation, ensuring that every data sample was included in the results. The XGBoost algorithm, a widely used tree-based ensemble method popular in biomedical research, is used for the machine learning model (Appendix 2). We use the area under the ROC curve (AUC) to calculate the performance for each dataset. Details of our testing environment are provided in Appendix 3.
@@ II-A Datasets
TABLE I: Data Characteristics. MS*=Metabolic Syndrome, BC*=Breast Cancer, HF*=Heart Failure, DR*=Diabetic Retinopathy, IBD*=Inflammatory Bowel Disease
@@
II-B Statistics
This study primarily involves grouping SHAP values based on feature values and analyzing these grouped SHAP values or comparing them across groups. The statistical experiments utilized in this study are as follows:
II-B1 Sign of SHAP Values
This experiment examines whether the SHAP values of a specific group significantly lean toward a particular sign (positive or negative), to decide whether a feature consistently contributes positively or negatively to the model prediction across certain value ranges.
One-sample t-test: This test checks whether the mean of SHAP values in a group is significantly different from zero. It assumes that the SHAP values are normally distributed. If the mean is significantly positive or negative, it provides evidence of the overall directionality of the SHAP values in that group.
Wilcoxon signed-rank test: As a non-parametric alternative, this test is used when the normality assumption is violated. It evaluates whether the median of SHAP values differs from zero, making it robust to outliers and skewed data distributions.
II-B2 SHAP Values Between Two Groups
This experiment assesses whether the differences in SHAP values between two groups are statistically significant.
Two-sample t-test: This parametric test compares the means of SHAP values between two independent groups. It assumes normality and equal variances across the groups. A significant result indicates that the mean SHAP value differs between the groups, suggesting that the feature’s contribution to the model prediction varies based on the group.
Mann-Whitney U test: When normality cannot be assumed, this non-parametric test is employed to compare the distributions of SHAP values between two groups. It is particularly useful when the data are ordinal or contain outliers, as it assesses whether one group tends to have higher or lower SHAP values than the other.
These tests are often applied to compare groups defined by binary feature values, such as comparing SHAP values for ”Male” vs. ”Female” in gender-related features.
II-B3 SHAP Values Between Multiple Groups
This experiment evaluates whether there are significant differences in SHAP values across more than two groups.
One-way ANOVA: This parametric test compares the means of SHAP values across multiple groups. It assumes that the data are normally distributed and have equal variances across groups. A significant result indicates that at least one group mean differs from the others.
Kruskal-Wallis test: When the assumptions of ANOVA are not met, this non-parametric alternative is used to compare the medians of SHAP values across groups. It evaluates whether the distributions of SHAP values differ significantly among the groups.
Post-hoc tests: If ANOVA or Kruskal-Wallis tests reveal significant differences, post-hoc tests (e.g., Tukey’s HSD for parametric data or Dunn’s test for non-parametric data) are conducted to identify which specific groups differ from each other. In this study, we utilized Tukey’s HSD as a post-hoc test.
II-B4 Regression Analysis
This experiment focuses on fitting SHAP values for continuous feature values to specific functions to evaluate the relationship between feature values and their contributions to model predictions.
Procedure: Feature values are set as the input variable (x), and SHAP values are the output (y). Each predefined function (e.g., linear, quadratic, exponential) is fitted to the data. The statistical significance of each function is evaluated based on the p-value of the coefficient of x, denoted as a.
Selection of the best function: If multiple functions are found to be statistically significant, the Root Mean Square Error (RMSE) is used to identify the function that best fits the data. This approach ensures that the selected function accurately represents the relationship between the feature and its SHAP values.
Applications: This analysis is valuable for identifying whether the impact of a feature is linear, non-linear, or follows a more complex pattern. For example, it can reveal whether an increase in a feature’s value consistently increases or decreases its contribution to the model, or if there are threshold effects.
II-C How to use the library
The description and installation guide for the CLE-SH package is accessible at: https://github.com/kyungjini/cle-sh. After installation, importing the package with import clesh will enable the analysis by executing clesh(features, SHAP values, label name) as in Figure2. Here, features indicates the dataframe containing the feature dataset used to train/test the machine learning model and calculate SHAP values. SHAP values are SHAP values which match to features. label name is required to generate more comprehensive literal sentences.
III Number of important features
@@
As statistical significance is a key factor in this study, we compare the average absolute SHAP values between features. As shown in Figure3.a, we calculate the statistical significance of difference between adjacently ranked features. When both groups of SHAP values satisfy the normal distribution test by the Shapiro Wilk test, a paired t-test is used (parametric); otherwise, the Wilcoxon rank sum test is used (non-parametric). By cutting the group of features by where statistical significance occurs, we determine the number of important features according to the number of cuts, as shown in Figure3.b. The minimal number of important features (candidate_num_min, default=10) and the maximal number of important features (candidate_num_max, default=20) are set by hyperparameters, as having too many important features can be ineffective and too few can be less informative. Between candidate_num_min and candidate_num_max, we select the number that shows the highest difference from the next cut, as shown in Figure3.c. Statistical significance is determined p-value¡0.05, which 0.05 can be changed by adjusting the hyperparameter p_feature_selection. In the clesh_result folder (see Figure2), the Figureshowing the number of important features according to the number of cuts, as in Figures 3.d and 3.e, will be saved.
IV Identification of feature data type
@@
V Univariate analysis
The overall process of univariate analysis is illustrated in Figure 4, highlighting the statistical tests and interpretation flow by feature type.
Figure 4: Flowchart of the univariate SHAP value analysis pipeline. This flow outlines the analysis procedure by feature type (binary, discrete, and continuous), showing how SHAP values are statistically processed and interpreted using hypothesis tests and curve fitting.
@@ V Univariate analysis
TABLE II: Univariate analysis in each dataset. Linear, Quadratic, and Sigmoid in the first column imply the number of features that fit best to each function with statistical significance.
@@
V-A Binary and Discrete Types
We provide box plots of SHAP values in each category, as shown in Figures 5.a and 5.b.
@@ V-A Binary and Discrete Types
V-B Continuous Types
SHAP values of continuous features are difficult to interpret using conventional statistics, as the relationship can be very complex to interpret or feature values can be too sparse. To balance the amount of information and utility, we first provide scatter plots (SHAP dependence plots without interaction analysis). Then, we apply three different functions for interpretation: linear (Eq. 1), quadratic (Eq. 2), and sigmoid (Eq. 3). By setting feature values as the input variable and SHAP values as the output, each function is fitted and the function is selected as in Section II-B 2) Regression Analysis. The selected function is displayed in the scatter plot. As shown in Table 2, the function type that fits the most is the sigmoid function, followed by quadratic and linear. This indicates that the relationship between feature values and SHAP values is usually non-linear, underscoring the need to analyze more than just the SHAP summary plot. An example of the linear function is illustrated in Figure 5.d, the quadratic function in Figure 5.e, and the sigmoid functions in Figures 5.f and 5.g.
$f(x)=ax+b\quad\text{(Eq. 1)}$
$f(x)=ax^{2}+bx+c\quad\text{(Eq. 2)}$
$f(x)=\frac{L}{1+\exp(-a(x-x_{0}))}+b\quad\text{(Eq. 3)}$
@@ V-B Continuous Types
VI Interaction analysis
Figure 6: Flowchart of the interaction analysis pipeline. The diagram illustrates how combinations of target and interaction feature types (binary, discrete, continuous) are handled, from feature pairing and grouping to significance testing and final reporting.
The overall process of interaction analysis is illustrated in Figure 6, showing how our pipeline systematically handles various combinations of target and interaction feature types. The analysis is performed for important features identified in Section III, referred to as target features. For each target, we define an interaction feature using the built-in SHAP utility function shap.utils.approximate interactions().
Interaction analysis is performed for important features in Section III, namely, target features. To define the interaction feature of each feature, we use the built-in function of SHAP library, shap.utils.approximate_interactions().
@@ VI Interaction analysis
VI-A Target: Binary/Discrete, Interaction: Binary/Discrete.
In each category of the interaction feature, the SHAP values of the target feature between different categories of the target feature are compared. The statistical test follows the method described in Section V-A. When a target feature is binary, a two-sample t-test (parametric) or Mann-Whitney U test (non-parametric) is applied. When a target feature is discrete with multiple categories, a one-way ANOVA (parametric) or Kruskal-Wallis test (non-parametric) is used, followed by Tukey’s HSD post-hoc analysis to identify statistically significant differences between categories.
VI-B Target: Binary/Discrete, Interaction: Continuous.
When the interaction feature is continuous, analyzing by its range can be difficult to interpret. To simplify, we group interaction values into two categories: those smaller than the average and those larger than the average. This allows us to treat the interaction feature as a binary feature and apply the analysis described in Subsection VI-A.
In Figure 7.a, an example of box plots in interaction analysis is illustrated (dataset: MS). The target feature ’Albuminuria’ shows a negative slope in its SHAP values, regardless of the interaction feature’s value (Triglycerides). However, in Figures 7.b and 7.c, we observe that the SHAP values, where ’Albuminuria’s feature value is 2, show a statistically significant difference from other SHAP values only when Triglycerides is above its average. This indicates that the impact of Albuminuria = 2 is more significant when triglycerides are above the average.
VI-C Target: Continuous, Interaction: Binary/Discrete.
SHAP values of the target feature are grouped by the category of the interaction feature. Each group of SHAP values is then regressed by the target feature values using the function type selected in Subsection V-B. Functions that are fitted with statistical significance will be displayed in the scatter plots as shown in Figure 7.d. If only one group (one category of the interaction feature) shows statistical significance for the fitting, there will be no comparison between categories of the interaction feature. If more than one group shows statistical significance for the fitting, a comparison is made to determine whether the pattern depends on the category of the interaction feature.
@@
As in Subsection VI-B, samples are divided into two groups based on whether the interaction value is below or above its average. The analysis follows the same procedure as in Subsection VI-C, treating the interaction feature as binary.
VII Discussion
This study aims to support researchers who seek to use SHAP analysis in a statistically valid and reliable manner by generating summary reports based on statistical significance. CLE-SH is designed to guide users toward more rigorous and reproducible usage of SHAP at the early and mid stage of analysis. To achieve this objective, we implemented a minimal yet essential pipeline that ensures statistical validity throughout the interpretability process. Section III describes the feature selection procedure, which identifies how many features should be primarily examined using SHAP values alone, without additional experiments. Section IV classifies each feature according to its variable type, an essential prerequisite for proper analysis. Section V conducts univariate analyses for individual features, and Section VI examines whether statistically significant interactions exist with the most interactive feature. The overall structure and example template of the automatically generated report are provided in Appendix 6, and complete PDF reports for each dataset are included in the source code repository. By integrating these modules into a unified workflow, CLE-SH bridges model explainability and statistical rigor. It enables users to obtain interpretable and statistically reliable explanations from SHAP values without manual configuration, thereby making SHAP analysis both rigorous and accessible.
The proposed library is intended to function primarily at the exploratory stage of data analysis rather than at the stage of final statistical confirmation. Because the appropriateness of a specific analytical approach depends on the characteristics of each dataset, it is neither practical nor desirable to predetermine a fixed analysis pipeline. Instead, CLE-SH aims to assist researchers in the early phase of investigation by summarizing statistically significant SHAP-based patterns, thereby helping them identify promising analytical directions. For this reason, the current version focuses on fundamental and essential functionalities, which naturally come with certain limitations but also provide many possibilities for future development.
As discussed in the main text, once the analysis moves into the interaction domain, it becomes inherently more complex and highly dependent on the specific data context. For continuous variables in particular, there may exist patterns beyond the functional forms considered in this study, and the presence or absence of interaction effects can vary depending on how variable boundaries are defined. In the current version, we analyzed only the single most prominent interaction for each dataset. However, interactions naturally occur at multiple levels and can be explored from various perspectives. For instance, Bayesian approaches could be adopted to model hierarchical dependencies among features, or large language models could be leveraged to generate more generalized statements about feature interactions.
As this library is intended to provide a summary report that can be used in the early exploratory and feature-screening stages of analysis, its applicability to large-scale medical data is an important consideration. Although CLE-SH has not yet been directly tested on very large clinical datasets, the statistical procedures implemented are relatively lightweight, and no computational issues are expected. However, since p-values are inherently dependent on sample size, very large datasets may yield inflated statistical significance even for negligible effect sizes. This suggests that further refinement could be made in selecting appropriate statistical techniques or by incorporating additional warning functions that alert users when significance might be driven primarily by sample size rather than true effect magnitude.
It is also worth noting that SHAP values can vary depending on the underlying predictive model. In this study, XGBoost was selected as the representative model for SHAP computation. Although CLE-SH can process SHAP values derived from any algorithm, tree ensemble models tend to align more naturally with the library’s statistical framework due to their discrete partitioning of feature spaces and the deterministic nature of TreeSHAP explanations [65, 66]. In contrast, neural network based models, such as multilayer perceptrons, generally require approximation-based SHAP methods, which are computationally more demanding and can be more sensitive to model parameters or initialization [67, 68]. Addressing these challenges and extending CLE-SH to support a wider range of model architectures will be an important direction for future work.
However, as the functionality of the library expands, it will gradually diverge from its original purpose as a lightweight screening tool for early-stage observation. Therefore, maintaining usability while incorporating more advanced analytical features will be a key challenge. Future development should focus on improving the user interface and accessibility—such as implementing a web-based platform and integrating large language models—to make the library more comprehensive and intuitive for diverse users.